OpenStudy (bananas1234):

Simplify √ x^100 y^16 would this be x^10 y^4?

Nnesha (nnesha):

you can change square root to exponent for example \[\huge\rm \sqrt[n]{x^m} = x^\frac{ m }{ n }\]

Nnesha (nnesha):

or another example \[\huge\rm \sqrt{x} =x^\frac{ 1 }{ 2 }\]

Nnesha (nnesha):

so change that square root to exponent

OpenStudy (bananas1234):

sorry i am still having trouble understanding.

Nnesha (nnesha):

here n is root for example in this problem n =2 bec there is square root and m is exponent which is under the square root